OpenStudy (anonymous):

in 2006, the number of Americans who were members of a health club was 42.7 milion people. This was an increase of 106% from the number of health club members in 1990. How many Americans were health club members in 1990? (Source:www.emaxhealth.com) Help me please

OpenStudy (dumbcow):

Let x be the num of members in 1990 42.7 = x+106% of x 106% = 1.06 42.7 = x+1.06x 42.7 = 2.06x x = 42.7/2.06 x = 20.73 million